OpenStudy (anonymous):

Simplify x - 1 ------ -------- x^2 + x - 6 x^2 - 6x + 8

OpenStudy (luigi0210):

is it set equal to zero?

OpenStudy (anonymous):

first solve x^2 + x - 6 and x^2 - 6x + 8

sam (.sam.):

\[\frac{x}{x^2+x-6}-\frac{1}{x^2-6 x+8}\] Factor the denominators

sam (.sam.):

\[\frac{x}{x^2+x-6}-\frac{1}{x^2-6 x+8} \\ \\ =\frac{x}{(x-2)(x+3)}-\frac{1}{(x-4)(x-2)} \] Then find the common denominator and stick those fractions together

OpenStudy (anonymous):

Ok thanks guys